alones (alones):

Which of the following points lies on the circle whose center is at the origin and whose radius is 10?

alones (alones):

\[(0,2\sqrt{5})\]

OpenStudy (aravindg):

x^2+y^2=r^2 is general eqn of circle with centre (0,0) and radius r.

OpenStudy (aravindg):

Use it here.

alones (alones):

Okay the formula u man this one? \[r=\sqrt{(x-0)^2+(y-0)^2}\]

OpenStudy (aravindg):

If the point satisfies the equation it is on the circle.

alones (alones):

*mean

OpenStudy (aravindg):

Yep same as yours.

alones (alones):

Oh lol mkay Oh so this one is false 100=(\(2\sqrt{5}\))\(^{2}\)+0

OpenStudy (aravindg):

Yeah false

alones (alones):

Same thing with this one 100=( \(\sqrt{10}\) )\(^{2}+0\)

alones (alones):

and the last choice for which i had is WAIT @AravindG IT SHOWS FALSEE ALSO 100=(\(5\sqrt{5}\))\(^{2}\)+\(\ (5\sqrt{5})^2\)

alones (alones):

O.mg. so all of them are false.. how's that possible >.>

OpenStudy (mayankdevnani):

what are the options?

alones (alones):

Mkay A. (√10, 0) B. (0, 2√5) C. (5√2, 5√2)

OpenStudy (mayankdevnani):

option C is correct

OpenStudy (mayankdevnani):

you made a typo in your solution

alones (alones):

Oh i jsut solved it.. and it doesn't mathces..

OpenStudy (mayankdevnani):

you took \[\large \bf 5\sqrt{5}~instead~of~5\sqrt{2}\]

alones (alones):

Oh shoot yea i took 5 ooh my bad xP

OpenStudy (mayankdevnani):

its okay !

jimthompson5910 (jim_thompson5910):

geogebra says that point C is on the curve. So you can use a visual tool like a graphing calculator to check